# GraphAU-Pain: Graph-based Action Unit Representation for Pain Intensity Estimation

### v1

@@ III-D Comparison to SOTA
III-E Ablation Analysis
@@
The ablation analysis in Table IV underscores the critical role of graph representation and GNN in the GraphAU-Pain model. Removing the graph representation (w/o graph rep.) reduces the mean F1-score from $66.2\%$ to $63.1\%$, mainly due to the performance drops in No Pain and Mild Pain, highlighting the value of graph modeling for capturing AU relationships. Similarly, removing the GNN layer (w/o GNN) causes a significant drop to $40.3\%$, emphasizing the importance of graph-based interactions in learning AU features. The simplest setup (Only ResNet), relying solely on CNNs, achieves the lowest mean F1-score of $35.2\%$, demonstrating that CNNs alone fail to effectively model AU correlations for pain estimation. These results affirm the superiority of graph-based learning methods for pain estimation.
TABLE IV: Ablation analysis for GraphAU-Pain.
@@
w/o graph rep. removes the graph representation from the model. w/o GNN removes the GNN layer. Only ResNet relies solely on CNNs. The highest value for each category and metric is bold; the second highest is underlined.
Model Metric No Pain Mild Obvious Mean
Full F1 93.1 51.2 54.3 66.2
Recall 92.9 52.5 51.0 65.4
Precision 93.4 50.0 58.1 67.2
w/o graph rep. F1 87.7 46.4 55.2 63.1
Recall 80.8 73.4 51.0 68.4
Precision 95.9 33.9 60.2 63.4
w/o GNN F1 86.4 30.7 4.0 40.3
Recall 79.9 43.2 10.2 44.4
Precision 93.9 23.8 2.5 40.1
Only ResNet F1 70.3 31.0 4.2 35.2
Recall 56.5 68.3 15.3 46.7
Precision 92.8 20.1 2.5 38.5
III-F Limitations
@@
While the proposed method advances pain estimation, it has key limitations. Firstly, since PSPI only captures facial expressions, it may not reflect true subjective pain [5]. Future research could focus on finding alternative pain indicators. Secondly, aligning UNBC with DISFA through UNBC+ removes three pain-related AUs and adds noise through predicted labels, potentially impacting performance. A promising direction is to design AU models specifically for pain-oriented datasets like UNBC. Lastly, although the AU occurrence-based ANFL module provides satisfactory representation, AU intensity-based approaches (e.g., GRN) could be explored, as AU intensity better relates to pain intensity.
IV CONCLUSION
@@
This paper presents GraphAU-Pain, a GNN-based model combining graph-based AU features with full-face representation for pain prediction. It surpasses the state-of-the-art methods while enabling AU-informed pain estimation for clinical transparency. GraphAU-Pain addresses challenges like limited data and class imbalance in the UNBC dataset. Key contributions include improved pain classification benchmarks, better interpretability through AU-based representations, and critical baselines for future AU-based pain estimation. Overall, this work demonstrates the potential of GNNs for accurate, clinically viable pain estimation solutions.
ETHICAL IMPACT STATEMENT
This study involved secondary analyses of an existing dataset. No human subject data are collected in this study. All experiments and analysis are conducted on a publicly available dataset that has been described and cited accordingly. Implementation code are open source, including preprocessing, model training, and post-hoc analysis, via github.com/ZW471/GraphAU-Pain.

### v2

@@ 4.4 Comparison to SOTA
LBP [48] 34.70 - - -
Head Analysis [49] 28.10 - - -
LPQ [48] 23.40 - - -
4.5 Ablation Analysis
@@
The ablation analysis in Table 4 underscores the critical role of graph representation and GNN in the GraphAU-Pain model. Removing the graph representation (w/o graph rep.) reduces the mean F1-score from $66.2\%$ to $63.1\%$, mainly due to the performance drops in No Pain and Mild Pain, highlighting the value of graph modeling for capturing AU relationships. Similarly, removing the GNN layer (w/o GNN) causes a significant drop to $40.3\%$, emphasizing the importance of graph-based interactions in learning AU features. The simplest setup (Only ResNet), relying solely on CNNs, achieves the lowest mean F1-score of $35.2\%$, demonstrating that CNNs alone fail to effectively model AU correlations for pain estimation. These results affirm the superiority of graph-based learning methods for pain estimation.
@@
Table 4: Ablation analysis for GraphAU-Pain. w/o graph rep. removes the graph representation from the model. w/o GNN removes the GNN layer. Only ResNet relies solely on CNNs. The highest value for each category and metric is bold; the second highest is underlined.
F1 Precision Recall
Model No Pain Mild Obvious Mean No Pain Mild Obvious Mean No Pain Mild Obvious Mean
Full 93.1 51.2 54.3 66.2 93.4 50.0 58.1 67.2 92.9 52.5 51.0 65.4
w/o graph rep. 87.7 46.4 55.2 63.1 95.9 33.9 60.2 63.4 80.8 73.4 51.0 68.4
w/o GNN 86.4 30.7 4.0 40.3 93.9 23.8 2.5 40.1 79.9 43.2 10.2 44.4
Only ResNet 70.3 31.0 4.2 35.2 92.8 20.1 2.5 38.5 56.5 68.3 15.3 46.7
4.6 Discussion and Future Work
GraphAU-Pain demonstrates significant potential for advancing automated pain assessment through several key contributions. By leveraging graph-based learning to model AU relationships, our approach achieves superior performance compared to existing methods, with an accuracy of 87.61% in the clinically meaningful three-category classification system. The model’s strong performance in detecting No Pain (93.10% F1-score) makes it particularly valuable for initial screening applications. Furthermore, the AU representation learning module provides a more interpretable framework for understanding how different facial expressions contribute to pain assessment. This could lead to more reliable and explainable automated pain monitoring systems in clinical settings, potentially reducing the burden on healthcare providers and improving patient care through continuous, objective pain assessment.
@@ 4.6 Discussion and Future Work
While the proposed method shows promising results in pain estimation, there remains room for improvement. Firstly, since PSPI only captures facial expressions, it may not reflect true subjective pain [50]. Future research could focus on finding alternative pain indicators. Secondly, aligning UNBC with DISFA through UNBC+ removes three pain-related AUs and adds noise through predicted labels, potentially impacting performance. A promising direction is to design AU occurrence prediction models specifically for pain-oriented datasets like UNBC. Lastly, while the AU occurrence-based representation learning module provides satisfactory representation, AU intensity-based approaches (e.g., GRN [21]) could also be explored since AU intensity better relates to pain intensity. However, this direction may require more complex models and additional training data to mitigate the impact of data imbalance.
5 Conclusion
@@
This paper presents GraphAU-Pain, a GNN-based model combining graph-based AU features with full-face representation for pain prediction. It surpasses the state-of-the-art methods while enabling AU-informed pain estimation for clinical transparency. GraphAU-Pain addresses challenges like limited data and class imbalance in the UNBC dataset through a novel transfer learning strategy. Key contributions include improved pain classification benchmarks, better interpretability through AU-based representations, and critical baselines for future AU-based pain estimation. Overall, this work demonstrates the potential of GNNs for accurate, clinically viable pain estimation solutions.
Acknowledgements.
Y. Liu’s work was supported in part by the Finnish Cultural Foundation for North Ostrobothnia Regional Fund under Grant 60231712, and in part by the Instrumentarium Foundation under Grant 240016.
Declaration on Generative AI
The author(s) have not employed any Generative AI tools.